1.    The vapor pressure of benzene is 100.0 mmHg at 26.1°C. Calculate the vapor pressure of a solution containing 24.6 g of camphor (C10H16O) dissolved in 98.5 g of benzene. (Camphor is a low-volatility solid.)


1
Expert's answer
2021-12-07T21:40:03-0500

P1= 100

P2= ?

V1= 98.5

V2= 98.5+ 24.6= 123.1

"V1P1= V2P2"

P2= 9850/123.1= 80.02mmHg